OpenStudy (anonymous):

solve 4/3+2/3q=-5/3-1/3q-4

OpenStudy (anonymous):

q=21

OpenStudy (anonymous):

I got different answer...

OpenStudy (anonymous):

4/3 +5/3 = 3 2/3q +1/3q= 1/q

OpenStudy (anonymous):

so, you have 1/q = -7 which gives you: q=-1/7

OpenStudy (anonymous):

what is the diference b/w our answers is u have taken 2/3q as 2 - 3q i took it as 2q -- 3

OpenStudy (anonymous):

you are right;) that as I saw it... I'm not sure what the original problem should be...?

OpenStudy (anonymous):

kkkk........